Hi, can you please help with the following? I need to solve these equations:

1. 2(h - 7) = -14

2. 1.6 + 4x = 8

3. (1/3)x + 5 = 7

4. (-2/3)x = 3/4



Answer by stanbon(48545)

1. 2(h - 7) = -14
divide both sides by 2 to get:
h-7=-14
Add 7 to both sides to get:
h=-7
----------------
2. 1.6 + 4x = 8
Subtract 1.6 from both sides to get:
4x=6.4
Divide both sides by 4 to get:
x=1.6
---------------
3. (1/3)x + 5 = 7
Subtract 5 from both sides to get:
(1/3)x=2
Multiply both sides by 3 to get:
x=6
--------------
4. (-2/3)x = 3/4
Multiply both sides by (-3/2) to get:
x=-9/8
------------
Cheers,
Stan H.